OpenStudy (anonymous):

You roll a standard number cube. Find P(number greater than 2) Answers A 5/6 B 1/3 C 1 D 2/3

OpenStudy (anonymous):

number cube = dice? if so then each side has a 1/6 chance of showing. for you the number has to be greater than 2 so, 3, 4, 5 or 6. each one has 1 in 6 chance of happening. so you have 4/6 = 2/3 chance of rolling greater than 2

OpenStudy (anonymous):

its (1/6) x 4 = 4/6